### v1

# TANQ: An open domain dataset of table answered questions

Abstract
Language models, potentially augmented with tool usage such as retrieval are becoming the go-to means of answering questions. Understanding and answering questions in real-world settings often requires retrieving information from different sources, processing and aggregating data to extract insights, and presenting complex findings in form of structured artifacts such as novel tables, charts, or infographics. In this paper, we introduce TANQ, [^1] Datasets at github.com/google-deepmind/tanq the first open domain question answering dataset where the answers require building tables from information across multiple sources. We release the full source attribution for every cell in the resulting table and benchmark state-of-the-art language models in open, oracle, and closed book setups. Our best-performing baseline, GPT4444 reaches an overall F1111 score of $29.1$, lagging behind human performance by $19.7$ points. We analyse baselines’ performance across different dataset attributes such as different skills required for this task, including multi-hop reasoning, math operations, and unit conversions. We further discuss common failures in model-generated answers, suggesting that TANQ is a complex task with many challenges ahead.
[…]
1 Introduction
Understanding and solving problems in real-world scenarios often requires reasoning across multiple documents and data modalities. This includes $(i)$ retrieving information from different sources, $(ii)$ processing and aggregating data to extract insights, and $(iii)$ presenting complex findings in a structured format, for example tables, charts or infographics, to communicate them to readers. Previous studies show that knowledge workers across domains, e.g. finance, science, and economics, spend around $20\%$ of their time searching and gathering information from different files into one document to extract insights and consequently answer information-seeking questions (Chui et al., 2012). As a result, one of most challenging tasks for workers is to aggregate data and turn it into insights. Tables as a structured representation of data are ubiquitous in real-world sources and are commonly used to communicate complex information. Hence, they can be the perfect modality to answer complex questions.
Figure 1: An example question in TANQ and its corresponding table answer. Supporting evidence from multiple pages in a Wikipedia snapshot is provided for each data point inside the table. We highlight the rationale inside each snippet in yellow. LLMs are evaluated with or without access to the evidence.
Dataset Open Domain Multi Doc Answer Type Document Type
Text Table Infobox
InfoTabs (Gupta et al., 2020) short text
FeTAQA (Nan et al., 2022) free form text
FinQA (Chen et al., 2021b) short text
TATQA Zhu et al. (2021) short text
MultiHiertt Zhao et al. (2022) numeric
OTTQA (Chen et al., 2021a) short text
NQ-TABLES (Herzig et al., 2021) short text
HybridQA Chen et al. (2020b) short text
MultiTabQA (Pal et al., 2023) table
TANQ table
Table 1: Comparison of TANQ to related (table) question-answering datasets.
Large language models (LLMs), commonly augmented with external tools such as retrieval, have become the go-to means for answering questions. State-of-the-art question-answering (QA) systems integrate LLMs in various ways, from decomposing complex queries to retrieving documents using external tools or generating context data from knowledge acquired during model training. However, their evaluation is mostly limited to simple datasets, e.g. TabFact (Chen et al., 2020a) or HotpotQA (Yang et al., 2018), whose questions can be answered by reasoning over a single table or text document and generating a short text sequence as answer. This limits the applicability of such systems to perform complex multi-step research explorations. Moreover, it differs from real-world needs where relevant information can be spread across documents and represented in different forms (e.g. text or tables). Moreover, more often than not, generating a short text as answer is not sufficient for more complex information-seeking questions.
In this paper, we investigate LLMs’ capabilities in reasoning over multiple data sources and formats (i.e. text, tables, infoboxes) to answer entity-centric questions and generate table answers as structured artifacts. To address these challenges, we introduce TANQ, an open-domain, multi-hop question-answering (QA) dataset. TANQ requires retrieving and aggregating data from multiple documents to compile and communicate answers as tables. To solve TANQ, models require different skills in addition to data retrieval such as filtering, maths, and name normalization. We create TANQ applying a five-step, automated data collection process. We use QAMPARI (Amouyal et al., 2022) as seed dataset and Wikidata as well as the Wikipedia corpus as data sources. For automated evaluation of different data collection and processing substeps, we use PaLM-$2$ (Anil et al., 2023b).
We evaluate several state-of-the-art LLMs on TANQ, including close, oracle and open book evaluation setups. Finally, we study model-generated answer tables and discuss common failure cases and challenges related to TANQ. Our evaluation of models across skills can further inform future tools and evaluation setups for LLMs to improve models for complex, information-seeking questions.
To summarise, our contributions are as follows:
We introduce TANQ, the first open-domain question-answering benchmark that requires building answers in form of tables from multiple information sources.
We benchmark state-of-the-art language models in oracle, open and closed book setups, reaching an overall F1111 score of $29.1$ with our best-performing (oracle) baseline.
We evaluate model performance across different dataset characteristics (e.g. reasoning skills and question types) and discuss challenges and common failure types.
Figure 2: TANQ creation pipeline consisting of five steps: 1. Extending QAMPARI questions; 2. Evidence extraction from Wikipedia; 3. Evidence evaluation and answer table extraction; 4. Question rephrasing; 5. Augmentation with additional skills. We include a running example at the bottom with a TANQ example at the end.
2 Related Work
Various benchmarks for QA have been released in recent years. Each one addresses different challenges related to the task. Table 1 provides an overview and comparison of benchmarks.
QA with text and/or table input.
A number of datasets use text and one or multiple tables for QA. While both text and tables have been considered as input modalities, the output of the datasets is mostly limited to short textual answers. HybridQA Chen et al. (2020b), for example, is a multi-hop QA dataset that requires reasoning over one table and multiple Wikipedia passages related to entities occurring in the table. HybridQA answers are short texts with location names being the most common answer types, followed by numbers, dates, and person names. Moreover, many QA benchmarks for reasoning over text and tabular context concentrate on the finance domain. For example, the MultiHiertt Zhao et al. (2022) benchmark is created from financial reports. Questions require reasoning over texts and multiple tables. The answers are short numerical values with a focus on numerical reasoning. Other financial QA benchmarks are FinQA Chen et al. (2021b) and TATQA Zhu et al. (2021) where the context is one table and minimum two paragraphs related to the table. TATQA answers are short texts consisting of either one or multiple text spans from context paragraphs/tables or are free-form answers. MultimodalQA is a multi-hop, open-domain QA dataset that takes one table and related images and text paragraphs as input with answers similar to previously described datasets.
Open-domain benchmarks.
Most of the earlier described datasets have context provided in form of text and/or tables, whereas open-domain QA datasets first require extracting the relevant context, before answering the given question. The majority of open-domain QA datasets, such as WikiQA Yang et al. (2015), TriviaQA Joshi et al. (2017), RobustQA Han et al. (2023), are limited to textual context. NQ Tables (Herzig et al., 2021) extends table-QA to an open-domain setting where first top-$k$ tables are retrieved from a given corpus. These tables are processed by a reader component for generating the correct short-text answer. Built on HybridQA, the Open Table-and-Text Question Answering (OTT-QA) benchmark (Chen et al., 2021a) extends this setting by requiring to extract both tables and texts given multi-hop questions.
QA with table answers.
The work closest to ours with respect to input and output modalities, is MultiTabQA Pal et al. (2023). MultiTabQA seeds on the Spider dataset, a text-to-SQL dataset containing SQL queries, database tables, and natural language translations of the queries. Pal et al. (2023) use table names occurring in SQL queries to extract the input tables and query Spider databases for answer table generation. While the dataset also generates answers as tables, it has certain limitations: $(i)$ the benchmark input is limited to tables; $(ii)$ the input and output tables are highly structured database tables which differ from real-world scenarios where tables occur in documents and websites in various formats; $(iii)$ the questions are limited to SQL queries formulated in natural language.
Question Type Example Question
1. Simple Which Belgian Grand Prix did Michael Schumacher win, and when did that happen?
2. Intersection For which movie did Chris Columbus receive credits as both director and writer and what was their composer, publication year, duration, and genre?
3. Composition Who choregraphed a work that was produced by the Royal Ballet? What was their date of birth, place of birth, occupation, and which awards did they receive?
Skill
1. No skill What filmmaker directed a movie written by Val Guest and what is their place of birth, date of birth, occupation, date of death, and place of death?
2. Filtering numeric Which film was directed and produced by Mel Brooks and what was their composer and duration? Filter the answer table for duration equal to or larger than 88 minutes.
3. Filtering time Which Italian footballer transferred to Pro Sesto in summer of 2020 and what is their date of birth, place of birth? Filter the answer table for date of birth equal to or after 1985.
4. Filtering entity What pieces of writing did Gregory Benford edit? What were their publication dates and publishers? Filter the answer table for publisher equal to Bantam Books.
5. Date-to-year conversion Who were the members of Black Sabbath and what was their year of birth, genre, instrument, and occupation?
6. Quantity conversion What work did Michael Mann write and direct and what was their publication date, duration in hours, genre, director of photography?
7. Time calculation Which governor of Connecticut died while in office? What was their place of birth, occupation, date of death, date of birth, political party and how many years did they live?
8. Approximation What are the townships in Harper County, Kansas and what is their population rounded to the nearest ten?
Table 2: An overview of question types in TANQ and additional skills we use for augmenting the TANQ questions for more complex reasoning beyond retrieval.
3 Building the TANQ benchmark
TANQ evaluates the capability to answer open domain, multi-hop questions by aggregating data and generating answer tables.
3.1 Task definition
A TANQ dataset instance is a triple $(q,t,D)$, consisting of an entity-centric question $q$, a table answer $t$, and a document set $D$ (see Figure 1). To answer the multi-hop question $q$, first multiple sub-answers are extracted from the document set $D$. The answer is generated in table form $t=\{t_{i,j}|i\leq n,j\leq m\}$ consisting of $n$ rows (i.e. one per extracted entities) and $m$ columns. The documents in $D$ provide supporting evidence for each cell of the answer table $t_{i,j}$. $D$ is either provided as input to models (oracle setting) or retrieved from the Wikipedia corpus as $D^{\prime}$ (open book) and can consist of texts, tables and infoboxes.
3.2 Preliminaries
QAMPARI. We use QAMPARI as seed dataset which is an open-domain QA dataset with lists of entities as answers. QAMPARI further includes Wikipedia text as supporting evidence for each entry of the answer list. Different to prior QA datasets with short textual answers, they align to natural questions which require a list of answers extracted from multiple sources. The dataset is semi-automatically created with Wikidata and Wikipedia as data sources. Moreover, Amouyal et al. (2022) ask human annotators to validate each QAMPARI question.
WikiData. Following Amouyal et al. (2022), we use Wikidata (WD) as a source for question generation. Wikidata (Erxleben et al., 2014) is a collaborative knowledge graph consisting of triples of entities and relations, i.e. $(e_{1},r,e_{2})$. Entities $e_{i}$ are values (e.g. $1990$) or items that represent a real-world concept, object, or a topic. [^2] https://en.wikipedia.org/wiki/Wikidata Relations are edges connecting entities, e.g. $(DonaldTrump,instanceOf,human)$. Whereas “Donald Trump” is the head entity and “human” the tail entity. We follow the notion for formal queries over Wikidata introduced by Amouyal et al. (2022) for QAMPARI. Applying a relation $r$ as query over WD items $e_{i}$ results in a set of (tail) entities: $[[r(e)]]=\{e_{i}|(e_{i},r,e)\in\text{WD}\}$.
Wikipedia. We extract supporting evidence in the form of sentences, tables and infoboxes (entity-tables at the top right corner of Wikipedia articles) from Wikipedia (WP).
Type Count Freq (%)
WD Item 57.6k 79.1
Time 11.9k 16.4
Numeric 3.2k 4.4
Text 46 0.1
Table 3: Extension entities ($1$st pipeline step).
Question Type Count Freq (%)
1. Simple 504 36.1
2. Composition 320 22.9
3. Intersection 571 40.9
Skill
1. No skill 385 27.6
2. Filtering numeric 85 6.1
3. Filtering time 236 16.9
4. Filtering entity 342 24.5
5. Date-to-year conversion 310 22.2
6. Quantity conversion 95 6.8
7. Time calculation 45 3.2
8. Approximation 94 6.7
Skills per question
1. No skill 385 27.6
2. One skill 843 60.4
3. Two skills 137 9.8
4. Three skills 30 2.2
Table 4: Question types and skills in TANQ. Simple denotes questions which require neither composition nor intersection. See Table 2 for exemplary questions.
3.3 TANQ benchmark pipeline
Figure 2 provides an overview of the TANQ pipeline outlining all steps for data collection, processing and evaluation. We use QAMPARI, Wikidata and Wikipedia as data sources, as well as PaLM-2 Anil et al. (2023b) for paraphrasing and validation. Applying a five-step process, we create various dataset entries with different question types and skills.
Step 1. Extending QAMPARI questions
Starting with the QAMPARI questions and answer lists of entities, we extend each question $q$ with additional WD relations using the answer entities $e_{a}$. QAMPARI questions are classified in either simple, composition (e.g. “Who directed movies screen-written by Steven Spielberg?”) or intersection (e.g. the example in Figure 1). We first query the WD knowledge graph to extract additional relations $r_{ext}$ linked to $e_{a}$, i.e. $r[e_{a}]=\{e_{ext}|(e_{a},r_{ext},e_{ext})\in\text{WD}\}$. Hence, each extension is a WD triple linking the QAMPARI answer $e_{a}$ (e.g. Hey Ram in Figure 1) through the relation $r_{ext}$ (i.e. composer) to a new extended entity $e_{ext}$, i.e. Ilaiyaraaja for relation composer and answer entity Hey Ram in Figure 1. We only select a relation which fulfils two conditions. First, it is part of a predefined relation set $R$, i.e. $r_{ext}\in R$. We manually specify $R$ based on the WD relation used to create QAMPARI questions. Second, the relation exists for all answers $e_{a}$ of question $q$. Given $n$ extension relations, which fulfil these conditions, we extend the question $q$ in a template-style fashion: “$[q]$ and what is their $[r_{ext1}],[r_{ext2}],[...]$ and $[r_{extn}]$.” For example, the question in Fig. 1, was generated based on the initial question “Which Indian movies were both directed and written by Kamal Haasan?” through extending with the WD relations composer and release year. Additionally, we extract all extension entities $e_{ext}$ of the extension triple.
Step 2. Evidence extraction from Wikipedia
Next, we collect for each extension triple ($e_{a}$, $r_{ext}$, $e_{ext}$) supporting evidence using Wikipedia as an evidence source. Hereby, we search for supporting text, tables and infoboxes in the WP articles of $e_{a}$ and $e_{ext}$. We apply simple heuristics and search for mentions of $e_{a}$ in the $e_{ext}$ article and vice versa. Moreover, we extend our queries with additional (heuristic-based) query words, for example considering different forms of representing numeric or time-related information occurring in queries.
Step 3. Evidence evaluation & answer table generation
To evaluate the correctness of the previously collected evidence texts, tables and infoboxes, we employ PaLM-2 as an evidence evaluator. We prompt the LLM to evaluate the extracted evidence in a natural language inference setting. For each extension triple $(e_{a}$, $r_{ext}$, $e_{ext})$, we construct template-based sentences $s$: “$\langle e_{a}\rangle\langle r_{ext}\rangle\langle e_{ext}\rangle$”, e.g. “Hey Ram composer Ilaiyaraaja” for row two in Figure 1. We query the LLM to label the sentence s as “supported”, “refuted”, “not enough information” based on the provided evidence in form of a sentence, table or infobox entry. We then only consider the triples supported by at least one piece of evidence. [^3] See Appendix A for the prompt template.
Figure 3: Baseline models evaluated on TANQ: closed book, oracle, and open book with augmented tools.
Answer table. Each extended relation corresponds to a column in the answer table (e.g. column “Composer” in Figure 1). Since the question in Figure 1 is extended with two additional relations (i.e. composer and release year), the resulting answer table has three columns. Hence, each cell in the answer table corresponds to a WD triple. For example, cell “Ilaiyaraaja” corresponds to triple ((((Hey Ram, composer, Ilaiyaraaja)))).
Step 4. Question rephrasing
This step increases the naturalness of template-based extension questions generated in Step 1. Similarly to the earlier step, we prompt a PaLM-2 model in a few-shot setting. To ensure the question meaning is preserved during rephrasing, we add structured annotations in parenthesis to questions with the name of each relation (e.g “Which Indian movies were both directed (directed_by) and written by (written_by) Kamal Haasan?”). We run up to 5 iterations of rephrasing and stop if all relations are present, or discard the question otherwise. [^4] See Appendix A for the prompt template.
Step 5. Augmenting with skills
Finally, to generate more challenging questions requiring further reasoning capabilities beyond retrieval, we extend TANQ questions by asking for realistic post-processing steps. Overall, we augment questions with the following additional skills: $(i)$ Filtering of answer table given a numeric, time, or attribute condition (rows $2.-4$ in Table 2); $(ii)$ Conversion of numbers, dates, locations and corresponding units (rows $5.-7$ in Table 2); $(iii)$ Calculation and introduction of an additional table column based on time attributes (e.g. “lifespan” in the 8. row in Table 2). $(iv)$ Approximating numbers to the nearest ten, hundred, etc., see last row (Table 2). We use up to three distinct skills for augmenting TANQ questions. Table 4 provides a breakdown of TANQ dataset samples across skills.
3.4 Dataset statistics and analysis
Overall, the generated TANQ dataset has $1395$ entries, $36.1\%$ of question type simple, $40.9\%$ intersection, and $22.9\%$ composition questions. Approximately $72.4\%$ dataset instances (i.e. $,01$ 010) require at least one additional skill to answer the question. See Table 4 for a breakdown of skills and question types in TANQ. TANQ questions have an average a length of $21$ tokens and require extracting information about three relations on average which are related to the answer entities (similarly to composer and release year in Figure 1). On average, TANQ answer tables have $6.7$ rows and four columns. Table 3 given an overview of extension entities used. We further plan to release a TANQ training set with approximately $42k$ samples.
Model simple composition intersection all
P R F1 P R F1 P R F1 P R F1
Oracle setting
PaLM-$2$ 16.8 17.7 15.4 13.3 14.1 11.6 20.1 23.8 19.5 17.3 19.4 16.2
Gemini 56.1 15.4 14.3 53.8 10.1 9.7 48.9 16.2 13.9 52.6 14.5 13.1
GPT4 44.3 29.7 29.0 34.8 20.1 22.7 35.7 23.8 22.0 45.7 29.4 29.1
Human 54.6 48.6 48.8
Closed book setting
PaLM-$2$ 17.2 17.1 13.6 17.0 10.3 10.4 26.4 22.0 21.1 20.9 17.5 15.9
Gemini 32.8 11.0 8.4 37.1 6.4 5.9 31.8 17.9 16.0 33.4 12.7 10.9
GPT4 32.5 20.7 17.4 27.3 12.5 12.6 45.1 26.0 24.3 37.3 21.5 19.6
Open book setting
Tool LM 13.4 12.5 9.7 21.5 14.1 15.1 19.4 19.2 15.7 17.3 15.8 13.2
Table 5: Baseline performance by question type. For all question types, we observe GPT4444 to outperform other baselines in oracle, closed, and open book settings, resulting in an overall F1111 score of $29.1$ for all question types.
Model Filter num Filter time Filter entity Date2Year Quantity conv Time calc Approx
P R F1 P R F1 P R F1 P R F1 P R F1 P R F1 P R F1
Oracle setting
PaLM-$2$ 16.3 19.6 16.5 15.8 20.2 16.0 14.3 18.6 14.0 19.8 20.8 19.2 14.3 11.3 10.7 17.0 14.8 14.9 11.7 14.8 10.0
Gemini 47.7 12.2 10.1 58.4 13.0 11.2 58.0 9.3 8.1 46.0 18.5 16.8 45.6 14.5 14.1 51.4 7.7 8.0 42.7 12.6 11.5
GPT4 24.0 20.1 21.1 42.6 26.4 27.0 27.7 30.0 28.0 44.7 23.7 23.3 34.6 14.7 15.2 56.7 28.0 29.5 24.3 7.7 7.7
Closed book setting
PaLM-$2$ 16.7 18.0 16.0 16.2 16.0 14.1 23.0 16.7 15.2 30.2 25.2 23.6 19.9 17.3 16.4 12.9 6.4 7.3 15.4 12.0 12.0
Gemini 28.2 14.5 11.1 23.7 13.9 11.3 37.3 10.2 8.2 36.0 18.0 16.0 19.0 15.0 13.6 51.5 1.6 1.8 33.0 9.8 9.3
GPT4 36.7 16.6 15.1 34.7 25.8 22.1 31.5 23.0 22.4 49.7 22.0 20.2 71.6 7.3 5.7 44.6 1.4 2.0 36.5 19.1 19.3
Open book setting
Tool LM 16.7 21.1 17.1 15.7 11.1 10.9 21.0 20.6 16.4 20.3 17.5 16.0 9.7 10.8 9.7 28.9 13.0 10.7 16.0 19.9 15.2
Table 6: Baseline performance by skills: Filtering with numerical/datetime/entity conditions, date-to-year (Date2Year), quantity conversion, time calculation, and approximation. In the oracle setting GPT4444 outperforms other baslines for most skills, while PaLM-$2$ performs better on some skills in closed book evaluation.
4 Baselines & Evaluation
We evaluate TANQ in $(i)$ closed book, $(ii)$ oracle, and $(iii)$ open book setups. We give an overview of the different approaches in Figure 3.
Closed book.
For all experiments, we use the PaLM-$2$ Unicorn model (Anil et al., 2023b), GPT4444 (Brown et al., 2020), [^5] https://platform.openai.com/docs/models/gpt-4-and-gpt-4-turbo and Gemini Pro (Anil et al., 2023a). We evaluate all baselines in a few-shot setting with three examples provided in the prompt (see top left prompt in Figure 3). The closed book setup evaluates LLMs’ capabilities in extracting relevant information acquired during training to answer TANQ questions.
Oracle.
In the oracle setup, we provide models source attribution for every cell of the answer table in form of oracle documents (i.e. text, tables, and infoboxes). Hence, the prompt consists of the TANQ question and multiple evidence sentences, infobox entries and/or tables (see bottom left prompt in Figure 3). To provide further context, we include the Wikipedia page title, for text evidence also sub-/section titles, from where the evidence was extracted. Moreover, we add randomly selected oracle documents from other questions with similar attributes. This makes the oracle setting more challenging and requires models to filter the correct evidence from all provided ones first.
Open book.
For the open book baseline, we extend the PaLM-$2$ Unicorn model with external tools for search and calculation. The Wikipedia-based search tools aim to mimic the human search approach on Wikipedia similar to Yao et al. (2023). The agent model can access Wikipedia information through three tools: $(1)$ WikiSearch(keywords): a keyword based Wikipedia search that returns the Wikipedia article most relevant to the provided keywords; $(2)$ FindEvidence(article, keywords): a article-specific search that returns matched sentences, tables and infobox entries given keywords and a Wikipedia article; $(3)$ GetIntro(article): returns the introduction section of a Wikipedia article. For calculations, we provide the model a Python engine as an external tool, $(4)$ Python(calculation). We design the agent model to decompose and solve the TANQ task in multiple sub-tasks, consisting of $(i)$ retrieving requested entities in an iterative manner (e.g. movies in Figure 1), $(ii)$ searching for entity-related information (e.g. release year), $(iii)$ post-processing information (e.g. filtering, calculation,, etc.), and $(iv)$ finally aggregating the information in form of a table. For some sub-tasks (e.g. entity retrieval), a separate sub-agent augmented with the required tools (e.g. WikiSearch) is spawned.
Human baseline.
We compare model performance against human performance based on $50$ answer tables generated by annotators. We provided the annotators the same input prompt as the models in the oracle setting (i.e. TANQ questions and evidence for each answer table cell).
Evaluation metrics
We evaluate answer tables using the RMS metrics introduced by (Liu et al., 2023). RMS views tables as unordered collections of mappings from row/colum headers to values. Hence, the metric is invariant to transpositions and column/row permutations. It allows small errors between tables keys/values of target and reference tables using the Normalized Levenstein Distance (Biten et al., 2019) with a threshold, which we set to $0.5$. The metrics returns both precision and recall scores.
5 Results & Discussion
In this section, we discuss the following research questions: $(1)$ Is TANQ a challenging dataset for state-of-the-art models? $(2)$ How well can models answer TANQ questions in a close book setting relying on knowledge acquired during training vs. extracted from provided context (oracle)? $(3)$ How well do tool-augmented models perform on TANQ? $(4)$ What are particular challenges related to question types, reasoning skills, complexity of questions, and the size of desired answer tables found in TANQ? $(5)$ What are common failure cases of evaluated models?
5.1 Question types
In Table 5, we compare the performance of all oracle, close, and open book baselines on TANQ. We find GPT4444 consistently outperforming other models with overall F1111 scores of $29.1$ and $19.6$ in oracle and closed book setting, respectively. While outperforming the other baselines, a large gap remains compared to our human baseline of $48.8$. While Gemini Pro (oracle) demonstrates higher precision compared to GPT4444 (oracle), there is still a meaningful gap for recall scores, indicating that Gemini successfully generates answer “subtables” containing only some of the requested rows. This aligns with our findings in Table 10, showing “missing entities” as the most prevalent failure type in Gemini (oracle) answers. In the oracle setup, most models perform best on simple questions, followed by composition and intersection. In the closed book setting, we observe higher performance on intersection questions than other types.
5.2 Reasoning skills
Model One skill Two skills Three skills
P R F1 P R F1 P R F1
Oracle setting
PaLM-$2$ 17.1 20.1 16.5 14.8 14.7 13.4 9.6 12.0 10.4
Gemini 54.3 13.9 12.3 42.2 12.6 11.4 58.0 8.9 8.1
GPT4 46.1 30.9 30.6 25.3 12.8 14.1 - - -
Closed book setting
PaLM-$2$ 23.2 18.7 17.2 18.1 15.5 14.6 19.9 20.9 18.8
Gemini 33.0 13.2 11.2 33.1 12.6 11.0 25.7 16.0 12.8
GPT4 35.6 22.7 20.8 59.8 11.5 10.6 - - -
Open book setting
Tool LM 20.1 17.5 15.3 15.5 15.1 12.5 14.8 22.1 16.8
Table 7: Baseline performance by number of skills used to augment TANQ questions.
We further augmented TANQ questions to generate more challenging questions requiring further skill sets in addition to search. Next, we discuss insights gained from baseline evaluation on each skill (see Table 6).
Model One Relation Three Relations Five Relations Ten Relations
P R F1 P R F1 P R F1 P R F1
Oracle setting
PaLM-$2$ 18.3 23.4 17.3 18.1 20.2 17.6 14.7 13.0 12.5 13.5 13.9 13.3
Gemini 62.2 15.6 13.2 53.6 15.0 14.0 38.7 13.1 13.3 20.8 6.6 7.8
GPT4 40.2 39.8 35.9 39.7 22.2 24.5 68.2 20.1 21.0 38.1 14.0 12.3
Closed book setting
PaLM-$2$ 20.2 19.0 15.4 20.5 17.0 15.4 15.8 12.3 12.3 20.5 17.7 18.5
Gemini 31.6 14.4 11.2 31.4 11.2 9.8 42.9 5.5 5.0 26.8 8.0 9.4
GPT4 32.7 32.9 27.9 29.2 15.9 15.1 43.8 5.3 6.5 25.5 0.2 0.3
Open book setting
Tool LM 16.0 19.2 14.3 15.6 14.4 12.3 14.0 7.7 6.2 3.8 2.0 2.4
Table 8: Baseline performance by number of requested relations in TANQ questions. This results in columns of answer tables as demonstrated in Figure 1. PaLM-$2$ outperforms other baselines as the number of relations increases.
PaLM-$2$ (oracle/close/open).
For PaLM-$2$, we observe that the model struggles particularly with questions requiring numeracy, i.e. approximation of numerical values, quantity conversion, calculations based on datetime attributes. This issue persists for the tool augmented model. While having the Python tool available for questions requiring calculations, the agent model often ignores this tool for calculations. Moreover, the tool augmented counterpart demonstrates better performance on filtering tables with numerical and entity conditions ($17.1$ and $16.4$ respectively) but there is still a meaningful gap for filtering based on datetime conditions.
Gemini Pro (oracle/close).
Comparing to other baselines, we find Gemini Pro struggling particularly with the filtering skill for both oracle and close book evaluation. Filtering the answer table based on an entity condition is seemingly more challenging for the model than the other filtering categories. Moreover, we observe in Table 6 a substantial decrease of the F1111 score for the skill “time calc” (i.e. calculations based on datetime attributes) when comparing oracle and close book evaluation and a substantial gap to other baselines in the oracle setting.
GPT4444 (oracle/close).
Across all skills except approximation, the GPT4444 (oracle) baseline outperforms its counterparts. In the oracle setup, GPT4444 struggles however with numerical filtering, quantity conversion, and approximation of numerical values. On the other hand, it outperforms other baselines in entity filtering and time calculation. In the closed book setting the model demonstrates decreased performance compared to other closed book baselines for multiple skills, such as numerical filtering, date-to-year conversion, and time calculation.
Number of skills.
Table 7 shows baseline performance by the number of skills required to generate the correct answer table. Naturally, we observe across baselines a decrease of performance as the number of skills increases and questions become more complex. However, for Gemini and PaLM-$2$ this gap is smaller than for GPT4444 (both oracle and closed book).
5.3 Relations in TANQ questions
Table 8 demonstrates baselines’ performance across the number of relations in the given question. For example, in Figure 1 two additional relations, i.e. composer and release year, are requested for the movies specified in the question. While all models show a performance decrease with an increasing number of relations, this gap is particularly significant for the tool augmented model where the F1111 score decrease by almost twelve points comparing questions requiring one relations vs. ten. Moreover, we observe for all baselines (oracle/close/open book) little performance drops comparing questions with one relation to those with three. The performance gaps increase as the number of relations increases further, i.e. from three to five and from five to ten.
Model Short Medium Long
P R F1 P R F1 P R F1
Oracle setting
PaLM-$2$ 12.8 20.2 12.9 14.7 19.9 14.0 18.8 19.1 17.4
Gemini 57.9 9.9 6.0 57.3 11.0 8.6 50.1 16.3 15.5
GPT4 37.6 40.2 38.4 39.2 34.4 34.0 49.1 26.7 26.5
Closed book setting
PaLM-$2$ 16.8 17.6 11.6 17.8 16.6 13.3 22.6 18.0 17.3
Gemini 41.5 8.1 3.5 31.9 12.1 8.5 34.3 13.1 12.2
GPT4 19.4 25.6 20.0 27.9 23.5 19.4 42.3 20.4 19.7
Open book setting
Tool LM 5.9 17.8 7.4 13.8 19.6 13.5 19.2 13.7 13.1
Table 9: Baseline performance by length of answer tables. Short tables have up to three rows, medium table up to six, and long table seven or more rows. For our best-performing baseline, GPT4444 (oracle), we observe a significant F1111 decrease with increasing table size.
5.4 Answer table length
Table 9 gives an overview of model performance across different answer table lengths: short answer tables (up to three rows), medium (up to six), and long tables (seven or more rows). A model seemingly struggling with longer tables is GPT4444 (oracle). The F1111 score decreases by $-4.4$ points comparing short to medium size tables and by further $-7.5$ points for long tables. It is apparent that the model struggles with extracting relevant information from the growing list of provided oracle documents as the table size increases. This is obviously not the case for close book evaluation as no oracle documents are given as part of the input.
5.5 Failures types
Failure Type Gemini PaLM-2 GPT4 Tool LM
Rel missing 37.3 33.3 0 3.9
No header 27.5 3.9 0 5.9
Filter issues 15.7 15.8 4.0 5.9
Halluc. relations 5.9 9.8 5.9 2.0
Halluc. other 0 15.7 0 7.8
Missing entity 80.4 58.8 27.5 78.4
Non-table 31.4 21.6 0 15.7
Partial answers 3.9 23.5 2.0 0
Wrong answer 14.9 7.8 9.8 13.7
[…]
We further study common failures cases of oracle baselines. The aim is to understand challenges related to TANQ when the model has access to necessary information, in form of oracle documents, for answer table generation.
Therefore, we manually evaluate a subset of model predictions to identify common failure cases which we organise in the following categories: 1.)1.)1 . ) Relation missing: At least one relation which corresponds to a column in the answer table is missing; 2.)2.)2 . ) No table header for the answer table is generated; 3.)3.)3 . ) Filtering (attribute, numeric, dates): Filtering conditions ignored or not applied correctly; 4.)4.)4 . ) Hallucinated relations: The answer table has columns which were not requested in the question; 5.)5.)5 . ) Hallucinations other other types of hallucination than additional relations or table format; 6.)6.)6 . ) Missing entity: entities, which correspond to rows in the table, are missing; 7.)7.)7 . ) Non-table Answers: Instead of generating a table as an answer, the model returns simple text or output in a different form than the table format provided in the prompt; 8.)8.)8 . ) Partial Answers: For a given entity and attribute only a partial answer is provided (e.g. only one university a person attended for relation “alma mater” although more exist); 9.)9.)9 . ) Wrong Answer (cell): Given answer (in a cell) is not correct.
Comparing open and oracle baselines, our observations show that the best performing model, w.r.t least failures, is GPT4444 (oracle). The only significant failure category we identify for GPT4444 are missing entities in generated answer tables. The most present issue we observe for Gemini Pro (oracle) are missing entities, missing relations, resulting in answer tables with a subset of columns, non-table outputs, and missing headers of generated tables. Similarly, multiple aspects pose a challenge for PaLM-$2$ (oracle), including hallucinations, missing entities/relations, and applying filtering conditions correctly. For the tool augmented model, one failure type is particularly prominent, namely missing entities that is followed by non-table outputs.
6 Conclusion
This paper introduces TANQ, the first open domain question answering dataset where the answers require building tables from information across multiple sources. To create TANQ, we design and apply an automated dataset pipeline using large language models and Wikipedia and Wikidata as knowledge sources. We further release for each cell of the answer tables source attribution in form of text, tabular or infobox proofs. We evaluate our dataset on state-of-the-art models in three different setting: oracle documents provided, closed, and open book setting. Our results and analysis suggest that TANQ is a complex task with many challenges ahead.
Appendix A TANQ Dataset
A.1 Prompt templates
Figure 4: Prompt used for evidence evaluation. We prompt a language model to evaluate the extracted evidence in a natural language inference setting. The LLM labels the input statements as “verifiable” or “not verifiable” based on evidence provided in form of a sentence, table or infobox.
Figure 5: Rephrasing questions increases the naturalness of template-based extension questions. To ensure the question meaning is preserved during rephrasing, we add structured annotations in parenthesis to questions with the name of each relation.
This section lists prompt templates we used for different steps of the TANQ dataset pipeline: evidence evaluation (Figure 4) and question rephrasing (Figure 5).